# Worst-Case Symbolic Constraints Analysis and Generalisation with Large Language Models

### v1

@@ 4.1 Worst-case Path Constraints Curation
For fine-tuning, we assembled a diverse dataset of 23,600 instances—comprising 17,700 training examples and 5,900 test examples—that spans a broad range of symbolic constraints, from simple arithmetic checks to complex logical compositions. This collection includes both synthetic programs and examples derived from real-world applications. Notably, a portion of the dataset comes from the Badger tool’s examples (Noller_2018), which capture a variety of behaviours representative of real-world applications.
@@
Independent from our fine-tuning dataset, we compiled a test set consisting of over 600 instances from 20 Java programs. Each instance includes ground-truth SMT2 constraints for several smaller values from a single Java program, designating a larger as the held-out target the model must synthesise (see Figure 2). By comparing the generated constraints against the held-out ground-truth, we directly measure the model’s ability to extrapolate the underlying symbolic rule rather than merely memorising observed patterns. These programs implement different algorithms, spanning array manipulation, searching, sorting, graph traversal, optimisation and recursion—see LABEL:appendix:benchmark_programs for the full list.
@@
To accommodate both modest and extensive leaps without incurring path explosion, we classify each instance into small, medium, and large tiers, each associated with distinct ranges of jump size, as shown and described in LABEL:appendix:prompts. As input size increases, synthesising the corresponding worst-case constraints requires more reasoning and results in longer outputs, often nearing or exceeding the model’s maximum generation capacity. While providing more examples in the prompt can improve output quality by offering richer context, it also encourages the model to generate longer reasoning chains, which consume a greater portion of the output budget. This introduces a critical trade-off: richer prompts can improve intermediate reasoning, but increase the risk that the model truncates before fully emitting the final constraint. To mitigate this, we impose an upper limit on prompt size to ensure adequate capacity is preserved for generation. When the full set of example constraints exceeds this limit, we reduce the prompt to three representative examples—the first, the median, and the last. If the reduced prompt still exceeds the threshold, we exclude the instance entirely. Although this may conservatively filter out potentially solvable cases, it serves as a practical safeguard: if three examples already exceed the output limit, the constraint for a larger is likely even longer and prone to truncation.
@@
After refining the benchmarking dataset through these constraints, we obtain a consolidated set of 671 instances. Each instance includes a prompt (featuring constraints at multiple smaller ) and asks for a constraint at a larger , alongside the corresponding solution. Restricting to at most 30 allows for practical verification using the Z3 SMT solver (z3-demoura), ensuring each instance can be definitively classified as correct or incorrect. See LABEL:appendix:benchmark_profile for a detailed analysis on the benchmark.
4.2 Reinforcement Learning from Solver-Guided Feedback
Our primary contribution is WARP-1.0-3B, a 3B parameter model fine-tuned to generate SMT2 constraints without any supervised data. We adopt the training framework proposed by sheng2024hybridflow. The post-training process comprises several key components:
@@ 4.2 Reinforcement Learning from Solver-Guided Feedback
We build upon Qwen2.5-3B (qwen2025qwen25technicalreport) and adapt it using a task-specific prompt designed to elicit SMT2 constraint generation. Illustrated in LABEL:appendix:prompt_template, the model is required to first reason through the problem before providing its final answer. This design choice, inspired by deepseekai2025deepseekr1incentivizingreasoningcapability, avoids imposing content-specific biases and allows the model’s intrinsic problem-solving capabilities to emerge naturally.
@@
Semantic Reward. For each predicted SMT2 constraint we call the Z3 solver to check logical equivalence with the ground-truth constraint . If Z3 proves , the episode receives a positive reward; otherwise it receives no reward. This binary signal is cheap to compute and correlates well with downstream verification performance described in subsection 4.3.
@@
reflecting our prioritisation of logical correctness over surface-level fluency. We intentionally leave the content of the <think> block unscored, focusing evaluation budget on verifiable constraint generalisation rather than natural language reasoning quality.
@@
4.3 Evaluation and Performance Metrics
Our evaluation protocol quantifies the performance of SMT2 constraint generation through automated, reproducible measures:
@@ 4.3 Evaluation and Performance Metrics
We define accuracy as the proportion of tasks for which the generated SMT2 constraint is both syntactically valid and semantically correct (i.e., successfully verified by Z3). By measuring solver-verified equivalence to the true worst-case execution condition, our metric eschews surrogate objectives such as BLEU or perplexity, which offer no guarantee of formal correctness.
5 Experiments
We assessed the ability of a range of large-scale and small-scale language models including our proposed WARP-1.0-3B to generalise symbolic constraints. Each model was evaluated over three independent trials, and mean accuracies are reported. The overall results are presented in Table 1.
@@ 5 Experiments
| Qwen/Qwen2.5-3B-Instruct | 11.18 | 9.99 | 8.20 | 9.79 (±1.49) |
| WARP-1.0-3B | 41.43 | 41.58 | 40.24 | 41.08 (±0.72) |
@@
5.1 Setup
Evaluation Prompt
We assess each model’s ability to produce complete, structured SMT2 constraints by using the uniform instruction set (see LABEL:appendix:evaluation-prompts). This instruction set directs the model to present its full reasoning between <think>…</think> tags and to return an exact, canonical SMT2 assertion between <answer>…</answer> tags, without any additional commentary or truncation. All evaluations are performed zero-shot, with no imposed limit on output length.
Baselines
We benchmark WARP-1.0-3B against nine open-source checkpoints (3B–32B) and several contemporaneous OpenAI API endpoints, selected to span a representative range of model scales within our computational budget.
5.2 Results
WARP-1.0-3B achieves an average zero-shot generalisation accuracy of on our worst-case symbolic-constraint benchmark, evaluated over three independent trials. This performance represents a more than fourfold improvement over the instruct-tuned Qwen2.5-3B-Instruct baseline () and nearly doubles the accuracy of DeepSeek-7B (). Furthermore, WARP-1.0-3B surpasses mid-scale models such as Falcon3-10B () and DeepSeek-14B (), while maintaining a trial-to-trial variance of less than percentage point, indicating strong stability and consistent solver alignment. Table 1 presents the complete zero-shot generalisation results for all models, including DeepSeek-R1-Distill-Qwen variants with 7B, 14B, and 32B parameters (hereafter DeepSeek-7B, DeepSeek-14B, and DeepSeek-32B, respectively).
On the other hand, the latest OpenAI thinking series model (o4-mini) shows a good performance against our benchmark. Nevertheless, the notable improvements obtained by our approach over the baseline show that our proposed solver-guided fine-tuning strategy is effective and can significantly boost a model’s capabilities in worst-case reasoning.
LABEL:fig:model-accuracy depicts each model’s average accuracy plotted against parameter count on a log scale. Compact baselines such as Qwen2.5-3B () and DeepSeek-7B () lie at the lower end, while Qwen2.5-7B (), Falcon3-10B (), and mid-sized open-source models like DeepSeek-14B () and Qwen2.5-14B () occupy the middle. Among the largest open models, DeepSeek-32B scores and Qwen2.5-32B scores , alongside proprietary model GPT-4o () and its successor GPT-4.1 (), whose sizes are not publicly disclosed but are believed to be comparable. Notably, WARP-3B achieves , outperforming its parameter-size peers and matching or surpassing several B models. WARP-1.0’s pronounced deviation from parameter-only scaling underscores the transformative effect of RLSGF, demonstrating that integrating formal SMT validation into the fine-tuning loop empowers compact LLMs to synthesise semantically precise worst-case path constraints with fidelity rivalling—and in some cases surpassing—much larger models.

### v2

@@ 4.2. Dataset Creation and Benchmarking
In order to accommodate the novel application of LLMs for the prediction of worst-case path constraints, a corpus of Java programs and the worst-case execution path constraints that follow have been collected and formatted to enable the development of generalisation and predictive capabilities in LLMs through reinforcement learning. For fine-tuning, we assembled a diverse dataset of 23,600 instances—comprising 17,700 training examples and 5,900 test examples—that span a broad range of symbolic constraints, from simple arithmetic checks to complex logical compositions. This collection includes both synthetic programs and examples derived from real-world applications. Notably, a portion of the dataset comes from SPF-WCA and the Badger tool’s examples (Noller_2018), which capture a variety of behaviours representative of real-world applications.
@@
Independent from our fine-tuning dataset, we compiled another test set consisting of over 600 instances from 20 Java programs. Each instance includes ground-truth SMT constraints for several smaller values from a single Java program, designating a larger as the held-out target the model must synthesise (see Figure 2). By comparing the generated constraints against the held-out ground-truth, we directly measure a model’s ability to extrapolate the underlying symbolic rule rather than merely memorising observed patterns. These programs implement different algorithms, spanning array manipulation, searching, sorting, graph traversal, optimisation and recursion.
@@
To accommodate both modest and extensive leaps without incurring path explosion, we classify each instance into small, medium, and large tiers, each associated with distinct ranges of jump size. As input size increases, synthesising the corresponding worst-case constraints requires more reasoning and results in longer outputs, often nearing or exceeding the model’s maximum generation capacity. While providing more examples in the prompt can improve output quality by offering richer context, it also encourages the model to generate longer reasoning chains, which consume a greater portion of the output budget. This introduces a critical trade-off: richer prompts can improve intermediate reasoning, but increase the risk that the model truncates before fully emitting the final constraint. To mitigate this, we impose an upper limit on prompt size to ensure adequate capacity is preserved for generation. When the full set of example constraints exceeds this limit, we reduce the prompt to three representative examples—the first, the median, and the last. If the reduced prompt still exceeds the threshold, we exclude the instance entirely. Although this may conservatively filter out potentially solvable cases, it serves as a practical safeguard: if three examples already exceed the output limit, the constraint for a larger is likely even longer and prone to truncation.
@@
After refining the benchmarking dataset through these constraints, we obtain a consolidated set of 671 instances. Each instance includes a prompt (featuring constraints at multiple smaller ) and asks for a constraint at a larger , alongside the corresponding solution. Restricting to at most 30 allows for practical verification using the Z3 SMT solver (z3-demoura), ensuring each instance can be definitively classified as correct or incorrect.
4.3. Reinforcement Learning from Solver-Guided Feedback
We also contribute WARP-1.0-3B, a 3B parameter model fine-tuned to generate SMT constraints for worst-case execution. For fine-tuning, we adapt the computationally-efficient framework proposed by sheng2024hybridflow with the following key components:
@@ 4.3. Reinforcement Learning from Solver-Guided Feedback
We build upon Qwen2.5-3B (qwen2025qwen25technicalreport) and adapt it using a task-specific prompt designed to elicit SMT constraint generation. As illustrated in 3, the model is required to first reason through the problem before providing its final answer. This design choice, inspired by deepseekai2025deepseekr1incentivizingreasoningcapability, avoids imposing content-specific biases and allows the model’s intrinsic problem-solving capabilities to emerge naturally.
@@
Semantic Reward. For each predicted SMT constraint we call the Z3 solver to check logical equivalence with the ground-truth constraint . If Z3 proves , the episode receives a positive reward; otherwise it receives no reward. This binary signal is cheap to compute and correlates well with downstream verification performance described in subsection 4.4.
@@
This reflects our prioritisation of logical correctness over surface-level fluency. We intentionally leave the content of the <think> block unscored, focusing evaluation budget on verifiable constraint generalisation rather than natural language reasoning quality.
@@
4.4. Evaluation and Performance Metrics
Our evaluation protocol quantifies the performance of SMT constraint generation through automated, reproducible measures:
@@ 4.4. Evaluation and Performance Metrics
We define accuracy as the proportion of tasks for which the generated SMT constraint is both syntactically valid and semantically correct (i.e., successfully verified by Z3). By measuring solver-verified equivalence to the true worst-case execution condition, our metric eschews surrogate objectives such as BLEU or perplexity, which offer no guarantee of formal correctness.
5. Experiments
We evaluate WARP, our neurosymbolic approach for incremental worst-case constraint reasoning, across a wide range of popular large language models to assess its general applicability. In addition, we present WARP-1.0-3B, a fine-tuned model trained with the WARP approach, which achieves the best performance in its size range and even surpasses several larger open-source and proprietary baselines. Together, these evaluations demonstrate both the broad effectiveness of WARP when applied to existing models and its ability to lift performance across scales, enabling models of different sizes to reach levels typically reserved for much larger ones. The evaluation of WARP is organised around three research questions:
- RQ1.:
How does incremental worst-case constraint reasoning enable LLMs to outperform vanilla prompting, particularly when extrapolating from small input sizes to larger ones?
- RQ2.:
To what extent does WARP’s performance improvement reflect genuine symbolic reasoning ability rather than domain-specific adaptation to SMT-LIB syntax?
- RQ3.:
How robust is incremental worst-case constraint reasoning across different model scales, varying extrapolation regimes, and program families representative of real software workloads, and what practical trade-offs emerge in terms of accuracy, scalability, and computational cost?
5.1. Evaluation Setup
Evaluation Prompt
We assess each model’s ability to produce complete, structured SMT constraints by using the uniform instruction set (see 4). This instruction set directs the model to present its full reasoning between <think>…</think> tags and to return an exact, canonical SMT assertion between <answer>…</answer> tags, without any additional commentary or truncation. All evaluations are performed zero-shot, with no imposed limit on output length.
Baselines
We benchmark WARP-1.0-3B against nine open-source checkpoints (3B–32B) and several contemporaneous OpenAI models, selected to span a representative range of model scales within our computational budget.
@@ 5.1. Evaluation Setup
| Qwen2.5-3B-Instruct | 11.18 | 9.99 | 8.20 | 9.79 (±1.49) |
@@
| GPT-5-2025-08-07 | 95.83 | 95.08 | 95.50 | 95.47 (±0.38) |
5.2. Overall Results
Our overall results in Table 2 first demonstrate that incremental worst-case constraint reasoning consistently equips LLMs to outperform vanilla prompting. While GPT-5 represents the strongest proprietary baseline, its performance under vanilla prompting falls short of all models evaluated with WARP when equipped with the incremental reasoning strategy. Specifically, GPT-5 achieves only 13% accuracy when prompted directly, but reaches 95.47% under WARP, underscoring that incremental worst-case constraint reasoning is essential for enabling reliable generalisation.
The fine-tuned model WARP-1.0-3B achieves an average zero-shot generalisation accuracy of on our worst-case symbolic-constraint benchmark, evaluated over three independent trials. This performance represents a more than fourfold improvement over the instruct-tuned Qwen2.5-3B-Instruct baseline () and nearly doubles the accuracy of DeepSeek-7B (). Furthermore, WARP-1.0-3B surpasses mid-scale models such as Falcon3-10B () and DeepSeek-14B (), while maintaining a trial-to-trial variance of less than percentage point, indicating strong stability and consistent solver alignment. Table 2 presents the zero-shot generalisation results for all models, including DeepSeek-R1-Distill-Qwen variants with 7B, 14B, and 32B parameters (hereafter DeepSeek-7B, DeepSeek-14B, and DeepSeek-32B, respectively).
The recently released GPT-5 achieves an average of on our benchmark when combined with the incremental worst-case constraint reasoning strategy, whereas applied directly it performs poorly (see Table 1). In parallel, our solver-aligned reinforcement learning framework (RLSGF) produces large gains at the small scale: WARP-1.0-3B improves from (instruct variant of base model) to . These results establish both the upper bound provided by frontier models and the relative improvement enabled by RLSGF at compact scales.
LABEL:fig:model-accuracy depicts each model’s average accuracy plotted against parameter count on a log scale. Compact baselines such as Qwen2.5-3B () and DeepSeek-7B () lie at the lower end, while Qwen2.5-7B (), Falcon3-10B (), and mid-sized open-source models like DeepSeek-14B () and Qwen2.5-14B () occupy the middle. Among the largest open models, DeepSeek-32B scores and Qwen2.5-32B scores , alongside proprietary model GPT-4o () and its successor GPT-4.1 (), whose sizes are not publicly disclosed but are believed to be comparable. Notably, WARP-1.0-3B outperforms its parameter-size peers and matches or surpasses several B models. WARP-1.0’s deviation from parameter-only scaling highlights the significant impact of RLSGF, demonstrating that integrating formal SMT validation into the fine-tuning loop enables compact LLMs to synthesise semantically precise worst-case path constraints, achieving fidelity comparable to much larger models.